OpenStudy (anonymous):

In this expression below shown below, what number does the letter n represent? 875 divided by 5=(n+70+5 ) divided by 5

OpenStudy (jlastino):

The two denominators cancel out so you are left with 875= n+70+5 875= n+75 875-75=n 800=n

OpenStudy (anonymous):

ur missing \[\div5\]

OpenStudy (jlastino):

You can take them out of the equation 875/5 = (n+70+5)/5 Multiply both sides by 5 5(875/5) = ((n+70+5)/5)5 875= n+70+5 875= n+75 875-75=n 800=n